OpenStudy (anonymous):

the infinite sequence of numbers (1, 1, 2, 3, 5, 8, 13, . . .), with each number from the third term on being the sum of the two previous numbers

OpenStudy (campbell_st):

you have Fibonacci numbers

OpenStudy (campbell_st):

the 1st 2 numbers are 1 and then each number is the sum of the previous 2 3rd term 1 + 1 = 2 4th term 1 + 2 = 3 5th term 2 + 3 = 5 etc

OpenStudy (anonymous):

Thanks

OpenStudy (campbell_st):

so in general, T = term and n = term number \[T_{n} = T_{n -1} + T_{n - 2}\] so the 3rd term is \[T_{3} = T_{2} + T_{1}\]